# A Subgrid Model of Electron-Scale Tokamak Turbulence

### v1

@@ V Conclusion/Discussion
A subgrid ETG model has been derived here which averages local electron-scale turbulence over intermediate scales in space and time to include in global ion-scale simulation. CBC simulations are carried out which show a clear scale separation in turbulent spectra in $k_{y}$. A kinetic form of the electron-scale electron heat flux is taken from local GENE simulation and added into global GEM ITG simulations using a simple Fick’s law diffusion model. Multiple ratios of ITG to ETG turbulent heat flux levels are considered, and the effects of increased electron temperature relaxation are described. A more accurate quasilinear heat flux model is constructed and compared against nonlinear ETG heat fluxes at multiple radial locations. Such a quasilinear model allows for the possibility of running a single nonlinear ETG simulation at the peak temperature gradient only, which can help expedite coupling of simulations at both scales.
Future work will consider the effects of the ETG potential on the ion-scale distribution of electrons, as described by (8). The diffusion coefficient, $D_{\textrm{ES}}$, can be found by following the motion of tracer electrons in the ETG field of a local ES simulation, and a theoretical model developed to capture radial properties of the diffusion. The ETG potential can then be added directly to local GEM ion-scale simulations, assuming periodicity on ion scales, and evolved in time to compare to the diffusive model.
@@
The effects of intermediate-scale zonal flows as shown in Fig. 2 might also be included as an extra global radial shear parameter. Theoretically this would require breaking the scale separation hypothesis as $\langle\delta f_{\textrm{ES}}\rangle_{\textrm{m}}\neq 0$ and $\langle\delta\textbf{E}_{\textrm{ES}}\rangle_{\textrm{m}}\neq 0$ when the spectra become multiscale in $k_{x}$. In multiscale scenarios where such intermediate-scale zonal flows are reported the ITG turbulence must be weakHolland, Howard, and Grierson (2017) or sub-critical,Howard et al. (2016) and so it would be prudent to first focus on including effects of ion-scale turbulence in electron-scale simulations per Ref. Hardman et al., 2019.
@@
Many topics exist for future research directions. These include adding the compressible effects and effects of ETG Reynolds stress which have been ignored here, considering any spectral transfer between scales, and proper coupling of ITG and ETG simulations to correctly capture effects of ion-scale turbulence in electron-scale simulations. Furthermore, it is important to understand when a scale separation hypothesis is valid, as electron transport spectra can broaden to become multiscale in the pedestal. Belli, Candy, and Sfiligoi (2022)
@@ Acknowledgements.
The authors would like to acknowledge Gabriele Merlo for helpful discussion about the GENE code and normalizations. This research was supported by the Frontiers in Leadership Gyrokinetic Simulation project, Scientific Discovery through Advanced Computing program, U.S. Department of Energy Contract #DE-SC0024425.
@@ Appendix A Normalizations
‘GN’ and ‘GM’ stand for GENE and GEM respectively. Eqs. 24 and 25 are used to convert the normalized velocity-space grid in GENE to SI values at multiple radial locations, and then to GEM normalized units. The velocity-space grid is recalculated for each particle depending on its radial location and particles are interpolated onto the new velocity-space grid to calculate $\hat{\Gamma}_{\textrm{ETG}}$. Approximately $95\%$ of GEM particles fall into the velocity-space grid at all radial locations, with the remaining particles outside the $v_{\parallel,e}$ domain ±$\pm 3v_{Te}$ set in GENE, where $v_{Te}=\sqrt{2T_{0e}(r)/m_{e}}$.
Furthermore, the magnitude of $\hat{\Gamma}_{0,\textrm{ETG}}=\langle(\textbf{v}_{E}\cdot\textbf{e}_{x})\delta f_{e}\rangle|_{r_{0}}$ needs to be converted to GEM normalized units. The quantities $v_{Ex}$ and $\delta f_{e}$ are normalized as follows
@@
where GENE includes a factor of $\rho^{*}=\rho_{D}/R$ scaling for the perturbed quantities $\delta\phi$ and $\delta f$. Radial variation in $v_{Ex}$ is ignored to use only the peak turbulence level. While GENE uses the radial basis vector $\textbf{e}_{x}=\nabla r$ by default, GEM uses the unit vector $\hat{\textbf{e}}_{x}=\nabla_{r}/|\nabla_{r}|$ for the radial dot product. For the circular geometry used this makes no difference, however this can be changed in GENE using the ‘norm_flux_projection’ flag if necessary.
Figure 9: (a) Comparison of source terms when integrating GENE data directly in velocity-space and when integrating by interpolating over particles in GEM. (b) Electron flux profiles when $\hat{\Gamma}_{0,\textrm{ETG}}$ integrated assuming only radial changes in geometry ($B(r,z),J(r,z)$).
Finally, $L_{\textrm{ref}}^{\textrm{GN}}=R$, while $L_{\textrm{ref}}^{\textrm{GM}}=\rho_{p}$, the proton gyroradius, so that the normalized second derivative (18) must be multiplied by a factor $(\rho_{p}/R)^{2}$, giving a total factor
@@
While this correctly accounts for differences in normalizations between the codes it does not yet give a radially constant $\hat{D}^{*}_{0,\textrm{ETG}}$ consistent with (16). This is because (28) depends radially on $n_{0e}(r)$. The other factor of $T_{0e}^{-3/2}(r)$ is cancelled by (24) and (25) when integrating over velocity-space. Further, when calculating the heat flux an extra $\frac{1}{2}mv^{2}$ factor will add another radial dependence on $T_{0e}(r)$. To see this, consider the GENE formula for the normalized heat fluxMerz (2008)
@@
which contains radial dependence on $\hat{p}_{0e}(r)=\hat{n}_{0e}(r)\hat{T}_{0e}(r)$. A caret (^) here indicates normalized variables.
The extra pressure factor is divided from (28) to retrieve an approximately radially-constant $\hat{D}^{*}_{0,\textrm{ETG}}$, and the source term is integrated on the original GENE velocity-space grid and converted to GEM normalization to compare to integration in GEM which sums over particle weights. The results of both integration methods are compared in Fig. 9 using the kinetic flux density of Fig. 4 and agree well. Note, some radial variation remains in $\hat{D}^{*}_{0,\textrm{ETG}}$ due to radial changes in $J(r,z)$ and $B_{0}(r,z)$ which are used when integrating (29). The effect of this is small and simply ignored. This remaining radial variation can be seen in Fig. 9, where the fluxes are calculated by integrating $\hat{\Gamma}_{0,\textrm{ETG}}$ in Fig. 4 using CBC equilibrium profiles.

### v3

@@ V Discussion
A subgrid ETG model has been derived here which averages local electron-scale turbulence over intermediate scales in space perpendicular to the field and in time to include in global ion-scale simulation. CBC simulations are carried out which show a clear scale separation in turbulent spectra in $k_{y}$. A kinetic form of the electron-scale electron heat flux is taken from local GENE simulation and added into global GEM ITG simulations using a simple Fick’s law diffusion model. Multiple ratios of ITG to ETG turbulent heat flux levels are considered, and the effects of increased electron temperature relaxation are described. A more accurate quasilinear heat flux model is constructed and compared against nonlinear ETG heat fluxes at multiple radial locations. Such a quasilinear model allows for the possibility of using a single nonlinear ETG simulation at the peak temperature gradient only, which can help expedite coupling of simulations at both scales.
@@
Future work will consider the effects of the ETG potential on the ion-scale distribution of electrons, as described by (8). The diffusion coefficient, $D_{\textrm{ES}}$, can be found by following the motion of tracer electrons in the ETG field of a local ES simulation, and a theoretical model developed to capture radial variation of the diffusion. The ETG streamer potential can also be added directly to local GEM ion-scale simulations, assuming periodicity on ion scales, and evolved in time to compare to the diffusive model. The effects of intermediate-scale zonal flows as shown in Fig. 2 might also be included as an extra global radial shear parameter. Theoretically this would require breaking the scale separation hypothesis as $\langle\delta f_{\textrm{ES}}\rangle_{\textrm{m}}\neq 0$ and $\langle\delta\textbf{E}_{\textrm{ES}}\rangle_{\textrm{m}}\neq 0$ when the spectra become multiscale in $k_{x}$. The formation of intermediate-scale zonal flows in multiscale simulation is dependent on the level of ITG turbulence,Holland, Howard, and Grierson (2017) and so it would be prudent to first focus on including effects of ion-scale turbulence in electron-scale simulations per Ref. Hardman et al., 2019.
@@
Many further topics exist for future research directions. These include adding the compressible effects and effects of ETG Reynolds stress which have been ignored here, considering any spectral transfer between scales, and proper coupling of ITG and ETG simulations to capture effects of ion-scale turbulence in electron-scale simulations. Furthermore, it is important to understand when a scale separation hypothesis is valid, as electron transport spectra can broaden to become multiscale in the pedestal. Belli, Candy, and Sfiligoi (2022)
@@ Acknowledgements.
The authors would like to acknowledge Gabriele Merlo for helpful discussion about the GENE code and normalizations. This research was supported by the Frontiers in Leadership Gyrokinetic Simulation project, Scientific Discovery through Advanced Computing program, U.S. Department of Energy Contract #DE-SC0024425. This research utilized resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility, through NERSC award FES-ERCAP0026751.
@@ Appendix A Normalizations
‘GN’ and ‘GM’ stand for GENE and GEM respectively. Eqs. 25 and 26 are used to convert the normalized velocity-space grid in GENE to SI values at multiple radial locations, and then to GEM normalized units. The velocity-space grid is recalculated for each particle depending on its radial location and particles are interpolated onto the new velocity-space grid to calculate $\hat{\Gamma}_{\textrm{ETG}}$. Approximately $95\%$ of GEM particles fall into the velocity-space grid at all radial locations, with the remaining particles outside the $v_{\parallel,e}$ domain ±$\pm 3v_{Te}$ set in GENE, where $v_{Te}=\sqrt{2T_{0e}(r)/m_{e}}$.
Furthermore, the magnitude of $\hat{\Gamma}_{0,\textrm{ETG}}$ needs to be converted to GEM normalized units. The quantities $v_{Ex}$ and $\delta f_{e}$ are normalized as follows
@@
where GENE includes a factor of $\rho^{*}=\rho_{D}/R$ scaling for the perturbations $\delta\phi$ and $\delta f$. Radial variation in $v_{Ex}$ is ignored to use only the peak turbulence level. While GENE uses the radial basis vector $\textbf{e}_{x}=\nabla r$ by default, GEM uses the unit vector $\hat{\textbf{e}}_{x}=\nabla_{r}/|\nabla_{r}|$ for the radial dot product. For the circular geometry used this makes no difference, however this can be changed in GENE using the ‘norm_flux_projection’ flag if necessary.
Figure 9: (a) Comparison of source terms when integrating GENE data directly in velocity-space and when integrating by interpolating over particles in GEM. (b) Electron flux profiles when $\hat{\Gamma}_{0,\textrm{ETG}}$ is integrated assuming only radial changes in geometry ($B(r,z),J(r,z)$).
Finally, $L_{\textrm{ref}}^{\textrm{GN}}=R$, while $L_{\textrm{ref}}^{\textrm{GM}}=\rho_{p}$, the proton gyroradius, so that the normalized second derivative, (19), must be multiplied by a factor $(\rho_{p}/R)^{2}$, giving a total factor
@@
While this correctly accounts for differences in normalizations between the codes it does not yet give a radially constant $\hat{D}^{*}_{0,\textrm{ETG}}$ consistent with (17). This is because (29) depends radially on $n_{0e}(r)$. The other factor of $T_{0e}^{-3/2}(r)$ is cancelled by (25) and (26) when integrating over velocity-space. Furthermore, when calculating the heat flux an extra $\frac{1}{2}mv^{2}$ factor will add another radial dependence on $T_{0e}(r)$. To see this, consider the GENE formula for the normalized heat fluxMerz (2008)
@@
which contains radial dependence on $\hat{p}_{0e}(r)=\hat{n}_{0e}(r)\hat{T}_{0e}(r)$. A caret (^) in (30) indicates normalized variables.
The extra pressure factor is divided from (29) to retrieve an approximately radially-constant $\hat{D}^{*}_{0,\textrm{ETG}}$. The source term is integrated on the original GENE velocity-space grid and converted to GEM normalization to compare to direct integration in GEM which sums over particle weights. The results of both integration methods are compared in Fig. 9 using the kinetic flux density of Fig. 4 and agree well. Note, some radial variation remains in $\hat{D}^{*}_{0,\textrm{ETG}}$ due to radial changes in $J(r,z)$ and $B_{0}(r,z)$ which are used when integrating (30). The effect of this is small and simply ignored. This remaining radial variation can be seen in Fig. 9, where the fluxes are calculated by integrating $\hat{\Gamma}_{0,\textrm{ETG}}$ in Fig. 4 using CBC equilibrium profiles.